OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 3 and . a) 2√2 b)√11 c) √13

OpenStudy (radar):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (radar):

\[(\sqrt{3})^{2}+(\sqrt{3})^{2}=c ^{2}\]

OpenStudy (radar):

\[3+3=c ^{2}=6\]

OpenStudy (radar):

\[\sqrt{c ^{2}}=\sqrt{6}\]

OpenStudy (radar):

\[c=\sqrt{6}=2.45\]